Select the option for "?" that continues the pattern in each question. 3, -6, 12, 4, 20, ?
Answer by Alex.33(110)

Just here to offer one way to understand this.
3-9=-6
-6+2*9=12
12-8=4
4+2*8=20
20-7=13
(and then 13+2*7=27)
In this way, ?=13. But just maybe. If 13 is in your list of options, perhaps I am right.